Pharoah Corporation has issued $1 million in preferred shares to investors with a 7.90 percent annual dividend rate on a par value of $100. Assuming the firm pays dividends indefinitely and the required rate is 10 percent, calculate the price of the preferred shares.
1
Expert's answer
2019-11-12T17:06:32-0500

The price of the preferred shares is:

P = D/r = 7.9/0.1 = $79.